# How to Evaluate Speech Translation with Source-Aware Neural MT Metrics

### v1

@@ 5.4 In-the-Wild Evaluation
These experiments show that, even under realistic conditions where ASR sources are generated fully automatically, our proposed alignment solution remains highly effective, yielding only a minimal degradation (2.0–2.5%) compared to the upper bound scores computed on manually segmented sources.
@@ 7 Limitations
On the Quality of ASR and MT Models
In Section 3.1, we discussed that MT models tend to exhibit higher quality than ASR models. In the following, we outline typical performance levels reported in the literature for both tasks, to contextualize our findings and clarify the expected reliability of the synthetic sources used in our experiments.
@@
For MT, as for ASR, performance scores vary substantially depending on domain, language pair, and, most importantly, the amount of available training data. Notably, the OPUS-MT Dashboard [^26] https://opus.nlpl.eu/dashboard/, visited on 29 Aug 2025 shows that high-resource language pairs achieve consistently high performance not only for closely related languages, such as English–French or English–German for which BLEU scores often exceed 40 and COMET scores 85-90, but also for linguistically distant ones, such as English–Chinese or English–Japanese. In contrast, substantial drops in quality are observed mainly for very low-resource languages, such as Yoruba or Wolof, where the scarcity of training data imposes severe constraints on model performance.
@@
The resulting values, reported in Table 15, allow to compare the resource demands of the ASR- and BT-based pipelines: since the generation of BT sources requires only the MT step, while the generation of ASR sources requires all four steps, the BT pipeline results 2–7 times faster than the ASR pipeline, depending on the level of hardware parallelism exploitation within the GPU via larger batchsizes. One of the factors driving this significant difference is the high cost of the segment boundary refinement Algorithm 1 (Table 15, id 4). It is important to note that the algorithm is not parallelizable by design, since shifting the right boundary of a segment changes the left boundary of the next segment (steps 7–9), preventing their concurrent processing. Ultimately, given the same available hardware and parallelism exploitation, BT sources are significantly more cost-effective to generate than ASR sources.
@@
The Two-Stage Re-Segmentation Algorithm at Work
Figure 6 illustrates how the two-stage re-segmentation algorithm works on a controlled example, where the objective is to realign the reference transcript, whose gold segmentation is disregarded, with the segments of the reference translation. The results of experiments conducted in this controlled mode are reported in Section 5.3.

### v2

@@ 5.4 In-the-Wild Evaluation
These experiments show that, even under realistic conditions where ASR sources are generated fully automatically, our proposed alignment solution remains highly effective, yielding only a minimal degradation (2.0-2.5%) compared to the upper bound scores computed on manually segmented sources.
@@ 7 Limitations
On the Quality of ASR and MT Models In Section 3.1, we discussed that MT models tend to exhibit higher quality than ASR models. In the following, we outline typical performance levels reported in the literature for both tasks, to contextualize our findings and clarify the expected reliability of the synthetic sources used in our experiments.
@@
For MT, as for ASR, performance scores vary substantially depending on domain, language pair, and, most importantly, the amount of available training data. Notably, the OPUS-MT Dashboard [^26] https://opus.nlpl.eu/dashboard/, visited on 29 Aug 2025 shows that high-resource language pairs achieve consistently high performance not only for closely related languages, such as English-French or English-German for which BLEU scores often exceed 40 and COMET scores 85-90, but also for linguistically distant ones, such as English-Chinese or English-Japanese. In contrast, substantial drops in quality are observed mainly for very low-resource languages, such as Yoruba or Wolof, where the scarcity of training data imposes severe constraints on model performance.
@@
The resulting values, reported in Table 15, allow to compare the resource demands of the ASR- and BT-based pipelines: since the generation of BT sources requires only the MT step, while the generation of ASR sources requires all four steps, the BT pipeline results 2-7 times faster than the ASR pipeline, depending on the level of hardware parallelism exploitation within the GPU via larger batchsizes. One of the factors driving this significant difference is the high cost of the segment boundary refinement Algorithm 1 (Table 15, id 4). It is important to note that the algorithm is not parallelizable by design, since shifting the right boundary of a segment changes the left boundary of the next segment (steps 7-9), preventing their concurrent processing. Ultimately, given the same available hardware and parallelism exploitation, BT sources are significantly more cost-effective to generate than ASR sources.
@@
The Two-Stage Re-Segmentation Algorithm at Work Figure 6 illustrates how the two-stage re-segmentation algorithm works on a controlled example, where the objective is to realign the reference transcript, whose gold segmentation is disregarded, with the segments of the reference translation. The results of experiments conducted in this controlled mode are reported in Section 5.3.